solve:
6/w = 8/w - 1/11
I am not doing well this week

Found 2 solutions by algebrahouse.com, edjones:
Answer by algebrahouse.com(1659)

6/w = 8/w - 1/11
11w[6/w = 8/w - 1/11] {multiplied entire equation by LCD, 11w, to clear fractions}
11w(6/w) = 11w(8/w) - 11w(1/11) {multiplied each term by 11w}
66 = 88 - w {cancelled and multiplied what's left}
-22 = -w {subtracted 88 from both sides}
w = 22 {divided both sides by -1}
www.algebrahouse.com

Answer by edjones(8007)

6/w = 8/w - 1/11
66=88-w multiply each side by 11w.
w=22
.
Ed